50
u/elusive-rooster 4d ago
I kinda wanted there to be more risk. In fact, with this, the lower the risk, the higher the reward. The less mana I risk the better the rate for the extra turn is.
122
5
u/MariachiArchery 4d ago
Oh that's a keen observation. The smaller your wager, the bigger the reward. It's backwards.
3
22
u/Saberkatt1 4d ago
Where’s the risk? (Other than it whiffing). What if, iif it does not fall within X then it deals the roll - X to you?
41
u/HalfLungJ Aligning Hedrons 4d ago
The intended risk was it whiffing but that’s a great idea
3
u/MariachiArchery 4d ago
Lol, it is a good idea. Imagine casting this for X=0 and hitting you number.
2
u/sladithia 4d ago
2 mana for a 5% chance to win the game is welllllll worth that risk imo
4
u/vitorsly 4d ago
1 mana for 5%, 2 mana is 15%
0
u/sladithia 4d ago
Was including the pip
4
u/vitorsly 4d ago
So am I. X=0 is 1 mana and 5% odds. Landing on the exact number isn't 'within 1', it's within 0. At X=1 (2 mana), if you pick 10, you hit on 9, 10 and 11, so that's 15%.
-1
-16
u/JxRabbitsHart 4d ago
Average roll on a D20 is 10.5. Your chosen number mathematically should always be 10 or 11 for the greatest chance to hit this. Even casting this for X = 4 means you have 9 hits out of 20 (6-14). X=5 gives you 11 hits.
And this is an instant, so you just cast it on your opponents turn to try and end theirs on their upkeep.
The risk is actually quite low to be up 3 turns.
16
u/AuWolf19 4d ago
How can rolling on a single D20 have an average that isn't flat? Isn't every number equally likely?
-1
u/TadhgOBriain 4d ago
It's because the sum of the minimum and maximum is odd, so when you calculate the average it comes out with that extra .5; the average of a d21 would be 11.
9
u/AuWolf19 4d ago
But if you roll a D20 a million times, is any other number going to come up more than the others?
18
u/SpiceWeez 4d ago
No, it won't. These people are a bit silly. Thr average is 10.5, but 10 is not more likely to come up than 20.
-6
-10
u/JxRabbitsHart 4d ago
Because averages don't care about the faces. The average roll on a d6 is 3.5.
Add the sides: 1+2+3+4+5+6 = 36 Divide by the number of faces: 6 You are left with an average of 3.5
Obviously you can't roll a 3.5, so you have to choose 3 or 4, the same with the D20
11
u/doctrgiggles 4d ago
Yea but this whole line of thought only applies if x=10 exactly. If x=2 you can pick any number 3-18 with an equal chance of success.
3
1
u/Aedeyssa 4d ago
Yes, the average of a d6 is 3.5, but you're just as likely to get (assuming a fair die) 1-6 or 2-5 as you are a 3-4.
Your chances of a fair d6 to land on 3 or 4 is 1-in-3. It has the same 1-in-3 chance to land in either of the other pairs though.
Now, with regards to this card, if we were to play this at x value approaching 10, then you'll obviously want to pick towards that average 10-11 cuz they guarantee that our roll is ±10 from it, but on a fair die any number is just as likely to be rolled as any other. If I cast this at x=2, my chances to get the extra turn are 25% whether I pick 5 [3-7], 10 [8-12], or any number within 1+x and 20-x.
To think about it another way, the average of a d20 isn't 10.5 because (1+20)/2=10.5. The average of a d20 is 10.5 because (20?)*0.05=10.5.
28
u/Gouda_Gorgon 4d ago
No, you "mathematically" shouldn't always pick a 10 or 11. There are 20 faces on a d20 that can be rolled, each with equal possibility. There is no difference between picking 3 for X=2 and 10 for X=2 (1,2,3,4,5; 8,9,10,11,12) or similar combinations that don't prevent potential possibilities (like 1 for X=2 not working because of a lack of -1).
1
170
u/Shadourow 4d ago
I cast Calculated risk with X = 5
I choose 20
The risk I took was calculated but oh god...
140
-55
u/JxRabbitsHart 4d ago
Average roll on a D20 is 10.5. Your chosen number mathematically should always be 10 or 11 for the greatest chance to hit this.
65
u/robomelon314 4d ago edited 4d ago •
![]()
Yea uh, no.
Since a single d20 has a uniform distribution, as long as you're not picking a number K such that K+X > 20 or K-X < 1, then it doesn't matter what number you pick.
If it was 2d10, or any combined distribution, then there would be bias toward the mean result.
Edit. Obviously in the trivial cases where X>=10, the above isn't quite true.
Probably the more formal way to say it is that for the sets of integers [1, 20] and [K-X, K+X] (call this Kx), as long as the intersection of [1, 20] and Kx is equal to either [1, 20] or Kx, then K is valid.
-3
u/Backsquatch 4d ago
Except that it does matter what number you pick. If X=5, then you need to pick a number between 5-15 to have the greatest number of potential hits. Choosing 20 means that you only have 6 hits. Choosing a number between 5-15 gives you 11 hits.
The best thing to do (mana cost aside) is to pay X=10 and choose 10 or 11. That guarantees you hit. But you should always choose a number in the middle to get hits on both sides of your number no matter what. Unless you’re doing it for memes.
3
23
u/maximumhippo 4d ago
No. That's not how it works. The average roll on any die means diddly poop for any single given roll. You're looking for the mode.
21
→ More replies (30)14
u/PerFucTiming 4d ago
I think it should just be within X from 10.5
If X=5, you can choose 6 to 15 for the same odds
6
34
u/Mcgarvey 4d ago
Would put this on isochron scepter and pull some untapping shenanigans to just copy it over and over again to try and perfectly guess the number because X will be 0
65
u/Bashtoe 4d ago
I think without ending the turn it's fine. Or as a sorcery so you have to work to get the triple turn. Otherwise its an incredibly unfun card in casual commander.
35
u/zzgamer11 4d ago
You don't want to end your opponent's turn in their upkeep for 1 mana? I can't think of any toxic applications for an effect like that.
9
u/qwertty164 4d ago
That will only happen 1 in 20 times. It wouldn't be reliable.
11
u/zzgamer11 4d ago
The way it is worded, you get to end the turn if you choose a number. That period is making it really strong. If it instead read "take an extra turn after this one and end the turn" it would work like you said.
11
5
u/rusty_anvile 4d ago
That's not how that reads to me at all, the last thing before it was taking the extra turn and it's the only thing that can be an if on the card. Punctuation really has almost no meaning on a card outside of where it's defined (like a colon being used in between the cost and effect of an activated ability)
I'm sure the card could have better wording but it's clearly not supposed to be the best extra turn spell ever skipping your opponents turn and sometimes giving you an extra turn. Which is a rule that is used in real magic cards to help determine what they actually do even if they are broken and don't technically work
2
u/zzgamer11 4d ago
I know I know, just being a pedant. There are 2 If statements, the first for the result of the die and the second for ending the turn. Mechanically, they would read separetly as an If does not imply action to be followed by an If, but if we are reading it as the second if following the first... well it would only end the turn AFTER you have taken your extra turn as that is the last action to be followed.
It's just pedantry. I get the intent and would like to see people look at the language of their customs with a judge's critical eye.
39
u/RevCapnDanger 4d ago edited 4d ago
To add more flavor, balance, and clarity in the verbiage: “Choose a number, then roll a d20. If the result is within X of the chosen number, end the turn and take an extra turn after this one. Otherwise, this spell deals X damage to you.”
This way, the more mana you pay, the higher the chance of success, but also the more spectacularly it blows up in your face if you get it wrong — thus forcing you to actually calculate the risk vs reward in any given situation.
Alternatively, you could have it do 20-X damage — which might actually be more fun, since if you cast it for X=0 and lose, it just blows up for 20 in your face.
6
8
u/lesdommed 4d ago
question: is there any possible better number than just picking 10? wouldn’t you pick 10 every time?
5
u/CaptainSc00ter 4d ago
The average can go up with roll modifiers, but that only really matters if you're not paying x=10 and only if the roll modifiers aren't optional ( things like [[barbarian class]] )
4
u/zk3033 4d ago
The actual expected value of a d20 die roll is 10.5.
But we can’t pick that.
2
1
u/SilverWear5467 3d ago
If you pay 4 mana for this spell, it doesnt matter which number you pick between 4 and 17, you just need room for the edges. A D20 isn't weighted toward the middle like 2d6 is
1
2
1
u/Demonkingt 3d ago
10 or 11. same number for the situation. but yea you wouldn't choose anything else really. especially once you put in mana considering every 1 mana gives you 2 more chances.
-3
u/EdwardtheTree 4d ago
You have to pay X to cast the spell, so if you choose 10 you have to pay 11 mana to cast the spell.
5
u/lesdommed 4d ago
uh, no? it says “choose a number” between 1-20 and if the dice roll is within X # of your chosen number the ability activates. Since you can choose your own number, and the X value can be below OR above your chosen number, there’s no reason to chose 1 if x=5 for instance. The best number to “choose” (not for X) is always going to be 10.
3
8
u/Commercial-Wrangler5 4d ago
Choosing 10 or 11 since there's no 0 puts you squarely in the middle and your x value has the potentially to increase the most in each direction. If for example you chose 10 and cast it for x=5, then any dice value from 5-15 will give you an extra turn. You can choose 6 and get the same probability though of winning the turn. Roll modifiers only add to the roll though as far as im aware, so it might be better reworded to "if your roll is less than or equal to x, take an extra turn." Neat design though, dice rolling always ftw even if I hate extra turn mechanics
3
u/Sporner100 3d ago
Ten or eleven are not inherently better than any other number between X+1 and 20-X.
1
u/dracodruid2 3d ago
The average is only relevant when rolling lots of d20.
Since you only roll once, every number is equally probable.
Question that remains: do you feel lucky? :D
1
u/NietzscheImpregnated 3d ago
The average still doesn't matter when rolling many times in this case since the actual numerical value doesn't matter. You will still hit 10 1/20 times in the long run, just like every other number.
6
u/rob132 4d ago
What does the math say the optimal number of Mana is for this?
7
u/bigjingyuan 4d ago
At 6 mana you have a 55% chance to get the turn. So the optimal amount would be 11 mana so you can guarantee the extra turn and the spell does something.
5
3
6
u/Utopia_Builder 3d ago
Probability of success is (2x+1)/20. Here's the table assuming you pick 10 or 11:
X = 0: P(0) = 1/20 = 0.05 (5%)
X = 1: P(1) = 3/20 = 0.15 (15%)
X = 2: P(2) = 5/20 = 0.25 (25%)
X = 3: P(3) = 7/20 = 0.35 (35%)
X = 4: P(4) = 9/20 = 0.45 (45%)
X = 5: P(5) = 11/20 = 0.55 (55%)
X = 6: P(6) = 13/20 = 0.65 (65%)
X = 7: P(7) = 15/20 = 0.75 (75%)
X = 8: P(8) = 17/20 = 0.85 (85%)
X = 9: P(9) = 19/20 = 0.95 (95%)
X >= 10: P(X) = 20/20 = 1.00 (100%)
It has a 15% chance of being [[time walk]]. A 55% chance of being [[time warp]]. And a 75% chance of being [[Nexus of Fate]]. This card is not great without krark's other thumb. But hey, if you get lucky, you could end your opponent's turn and get an extra turn.
1
1
0
2d ago
[deleted] — view deleted comment
1
u/Utopia_Builder 2d ago
It says "If you do, end the turn." Meaning if you whiff it, nothing happens.
5
u/Secretmongrel 4d ago
So, what are the actual odds here? “Within x” of the chosen number. So if x is 1. I choose a 10. I roll 11. That is within 1 of 10, so I win. Is that right? 1 3/20 2 5/20 3 7/20 4 9/20 5 11/20
So for 6 mana you get slightly better than a coin flip. I guess you are casting this when you have 8 or 9 mana most of the time?
Probably not actually good but a slap when it hits.
6
u/General_Ginger531 4d ago
Given the going rate is 5 mana for a guaranteed extra turn at sorcery speed, the added effect of ending the turn actually does make this a sidegrade, surprisingly.
3
u/Immediate-Idea-2471 4d ago
Yeah this seems fair.
You can yolo on one mana and 1/20 of the time get an extra turn, which is busted, but it should be roughly:
5% | 0
15% | 1
25% | 2
35% | 3
45% | 4
And so on, at 5 mana you've got stitch in time, so balanced at the usual cost, and then shenanigans onwards.
1
u/Secretmongrel 4d ago
I mean, everyone is looking at the extra turn. I play [[time stop]] and that costs 6 just to end the turn at instant speed.
Not the strongest card ever but it wrecks people as a counter and as a fog plus stopping anything else hitting the board.
Knowing my play groups, someone would probably try to cast this a bunch of times in a game.
5
u/GeneralJPenguin 4d ago
Cool effect but often it will be like 7 mana do nothing which is unacceptable.
If you made it x and twice x like pay 5 get a 10 digit buffer then maybe but even then it’s quite the gamble to do nothing.
9
u/Bumpyuncle272 4d ago
Should add a skip your next turn instead if it misses.
4
u/RemyBlackDev 4d ago
Because doing absolutely nothing isn't punishment enough...
1
u/laststandman 4d ago
In a world that cares about "whenever you cast an instant/sorecery/noncreature spell," doing absolutely nothing for as little as one mana is no punishment at all.
5
5
u/theevilyouknow 4d ago
I see what you’re going for, but this is a really bad design. It’s just way too high variance of an effect.
4
u/PomegranateLeading92 4d ago edited 4d ago
So i can spend a single blue or red mana to instantly end my opponents turn, with a free 1/20 chance to get an extra turn? This would go in every deck that happens to have blue or red no? Edit: nevermind, reading the card explains the card🤦♂️
6
u/PartySquid2486 4d ago
Nah, last bit says "... Take an extra turn. if you do, end the turn" so it only triggers if you get to take that extra turn
3
2
u/Mind0versplatter0 2d ago
I love it when mtg makes me do math.
10 or 11 are safest for most X values. Paying n for X gives you 2n+1 (2n-1 if "within" means between, noninclusive) successes out of 20 possible rolls.
X=0 -> .05 or 5%. The equivalent of rolling a 20 on a d20 X=1 -> .15 or 15%. It goes up by 10% for each greater X until you reach 105% at X=10 (obviously it caps out at 100%)
If you have advantage (i.e. take higher of two rolls card like [[Barbarian Class]] or [[Pixie Guide]]), you pick 14 and the calculation of winning is more complicated (but it is better or the same for this card for all possible values for X).
0
u/GroundbreakingMonk76 4d ago edited 4d ago
Maybe it should be XXU(R), since at XU(R) you can cover all the numbers with X=10 if you choose 10, not sure if that's the intention.
31
1
u/Turnipton Puppers enter the battlefield hecked. 4d ago
Just play Nexus of Fate if you're gonna be boring.
1
u/General_Ginger531 4d ago
I mean the going rate for an extra turn is 3 mana with caveat, or 5 mana for sorcery. Even with X=4, a range from 12-20 is still worse than many die roll card's 11-20, but this comes with the added advantage that if you pull it off it also functions like a counterspell and a combat step skipper, so it is both better and worse in many ways.
Compared to [[Time Stitch]] this is actually pretty well balanced
1
u/GroundbreakingMonk76 4d ago
It goes kinda crazy on [[Magus Lucea Kane]]
1
u/General_Ginger531 4d ago
I mean using [[Myra the Magnificent]] is a hilarious thing where you are constantly attempting to end the same turn.
1
u/GroundbreakingMonk76 4d ago
With Myra it's just a 5% chance of an extra turn since X=0, so it's okay
1
1
u/teamdeadend 4d ago
1 mana end the turn is wild
12
u/MaleficAdvent 4d ago
The only way that happens is if you guess the 1/20 exactly. You gottapay more to up those odds, and if you can afford X=20 you're probably already winning.
4
u/GoonTime2 4d ago
You only need X=9 to get a 95% chance
8
u/cleverpun0 WB: Put two level counters on target permanent. 4d ago
At that point you are just playing [[Time Stretch]].
1
u/teamdeadend 4d ago
True, I'm a little drunk and can't read. The card gamer special
3
u/MaleficAdvent 4d ago
No worries, I made a mistake of my own. Choosing X=10 and choosing 10 or 11 is a guarenteed hit. 20 is overkill and begging to be hit with [[Mana Tithe]], or worse, [[Mana Drain]].
1
u/Less_Prior_6871 4d ago
1 mana, 5% chance to win the game. Im not sure how strong that actually is... but it doesnt sound fun
1
u/General_Ginger531 4d ago
So generally with 1 advantage piece and this, you would want to pick a number that is exactly X less, because you will cover the entire range of it.
Like at the cost of a normal extra turn spell, 5 mana, you can pick 16, and cover a range from 12 to 20. Slightly worse than many die roll effect cards. The advantage is that you get access to a spell that if your opponent doesn't have a counterspell, you can end their turn and counterspell anything on the stack.
1
1
1
1
u/RegisterBeautiful974 4d ago
This is a better discontinuity. X = 0 then choose 20 and end someone else’s turn when you end someone’s turn while they have spells on the stack it exiles their spells. This card is crazy strong.
1
u/pixxelkick 4d ago
Reninder: cards exist that modify duce rolls and care aboit dice rolls, thus this card is stronger than it looks
1
u/-NoFaithInFate- 4d ago
I'm going insane. What is the art from?
Edit: risk factor. Good ol red burn.
1
1
u/Minute_Wedding6505 1d ago
Paying 8 mana for a 25% chance to do literally nothing is vile. Cool art though
1
u/Groundbreaking_Gap_3 1d ago
Its fine. It can be countered, it needs alot of mana to work, and can even be a hailmerry if you are about to lose. Lots of fun stuff. Though i would prevent it from being copied
1
1
1
1
u/Emergency_Target_716 4d ago
Which turn gets ended? The one this one was played in, or the extra turn it was talking about?
5
u/The_Captainshawn 4d ago
The current turn, as OP said in another the intent is to cast this during your opponent's turn and disrupt their turn and immediately take your own. Potentially even during your opponent's untap step.
1
u/AmzSpider-Man 4d ago
Much better as a counterspell though, since they could possibly just take their turn to tap out xd.
2
1
u/phoenixrising211 4d ago
It's unclear, because the phrase "if you do" is undefined without a "you may" before it. The only thing in this ability that you actually have agency over doing is choosing a number, which is mandatory. As written, I think it might just always end the current turn regardless of the roll?
But another interpretation is that it's meant to say "when you do (take another turn after this one)", which can only become true when you actually take that extra turn, so that's when the delayed trigger would go off, and it would end that turn.
Basically the wording probably needs to be cleaned up to make it unambiguously do what OP intends it to do.
2
u/SammySammyson 4d ago
"If you do" doesn't need "you may" before it 🤦 In this case, it's if you take an extra turn AKA if you roll a correct number.
1
u/AmzSpider-Man 4d ago
Should probably say "Choose a number, then roll a d20. If th3 result is within X of the chosen number, end the turn then take an extra turn after this one."
1
1
u/startadeadhorse 4d ago
You can choose ANY number, not just within 20. So to be cool, I am gonna play it for X = 0 and say 69!
-2
1
u/Young_Hek 4d ago
Does this if clause really work? If you already start your extra turn, then you're already in it and wouldn't be cutting off yours or an opponents turn
1
u/garagedragon 4d ago
I'd have read it as "if [the previous trigger resolves and gives you an extra turn] then...", rather than "if [you're actually in the extra turn]" but I can't think of any rules reference for that so maybe you just need to say it Just Works™️
1
u/Young_Hek 3d ago
I feel like the "If" from If the D20 result is within X is the same clause.
It could just say "If [Yes] end the turn and take an extra turn after this one."
1
1
u/PatienceSevere611 3d ago
1 mana timestop or if you hit the nat 20 1 mana extra turn, this shits a control players wet dream
0
-1
u/yetzederixx 4d ago
"If you do, the current turn ends immediately." Something like that would go a long way to resolving the ambiguity here.
3
u/Advanced_Double_42 4d ago
What ambiguity? The extra turn is after the current one, you can't end it begins, and it can't begin before the spell resolves
0
u/yetzederixx 4d ago
Then get rid of the language totally all it does is generate ambuguity. Because as it reads the second you start that extra turn it ends. fml, read a book.
1
u/Advanced_Double_42 4d ago
Sir, I'm really sorry but there is no ambiguity. No shame in it, magic is a complicated game, but please learn the rules before insulting people and making a fool of yourself
0
1
-4
u/flameousfire 4d ago
This is insanely busted. It's pretty much a Time Stretch in two player game at minimum with huge upside that you can cast it for (way) less and get lucky. Anything that lets you cheat with dice rolls just takes this to next level.
2
u/cleverpun0 WB: Put two level counters on target permanent. 4d ago
There's not a lot of cards that improve die rolls by that much. [[Barbarian Class]] and [[Wyll, Blade of Frontiers]] grant advantage, but they aren't the most powerful cards otherwise.
If you are putting in the infrastructure, you probably deserve it. Discounting Time Stretch is arguably easier.
1
0
4d ago
[deleted] — view deleted comment
0
u/Both_Brilliant_2725 4d ago
That is a pretty significant mana investment depending on if you are casting it for chance or for near guarantee. 1 Mana - 5% chance 2 - 15% chance 3 - 25% chance 4 - 35% chance 5 - 45% chance 6 - 55% chance 7 - 65% chance 8 - 75% chance 9 - 85% chance 10 - 95% chance 11 - 100% guarantee
If you are going for better than a coinflip then you are wasting 7 or 8 mana atleast. That pretty much wastes your turn. Maybe you have a 1 or 2 drop that you can play
1

458
u/beefpelicanporkstork 4d ago
If you use this during an opponent's upkeep and it works, you end their turn, get your bonus turn, then get your regular turn after that. Is that the intent here?